# Finding Dori [Uncaptioned image]: Memorization in Text-to-Image Diffusion Models Is Less Local Than Assumed

### v1

@@ E.3 Starting Embedding Optimization from Random Embeddings
Setting Adv. Steps Memorization Type $\boldsymbol{\downarrow}\bm{\text{SSCD}_{\text{Orig}}}$ $\boldsymbol{\downarrow}\bm{\text{SSCD}_{\text{Gen}}}$ $\boldsymbol{\downarrow}\bm{D_{\text{SSCD}}}$ $\boldsymbol{\uparrow}\bm{A_{\text{CLIP}}}$
NeMo + [Uncaptioned image] $1$ Verbatim $0.07\pm 0.02$ $0.06\pm 0.02$ $0.31\pm 0.06$ $0.21\pm 0.02$
Template $0.10\pm 0.03$ $0.10\pm 0.03$ $0.26\pm 0.05$ $0.28\pm 0.02$
NeMo + [Uncaptioned image] $10$ Verbatim $0.81\pm 0.06$ $0.89\pm 0.05$ $0.99\pm 0.01$ $0.32\pm 0.01$
Template $0.42\pm 0.13$ $0.21\pm 0.15$ $0.72\pm 0.14$ $0.32\pm 0.02$
NeMo + [Uncaptioned image] $25$ Verbatim $0.88\pm 0.04$ $0.95\pm 0.03$ $1.00\pm 0.00$ $0.32\pm 0.01$
Template $0.50\pm 0.13$ $0.20\pm 0.15$ $0.75\pm 0.11$ $0.32\pm 0.02$
NeMo + [Uncaptioned image] $50$ Verbatim $0.91\pm 0.03$ $0.97\pm 0.02$ $1.00\pm 0.00$ $0.33\pm 0.02$
Template $0.55\pm 0.12$ $0.18\pm 0.12$ $0.79\pm 0.10$ $0.32\pm 0.02$
NeMo + [Uncaptioned image] $100$ Verbatim $0.93\pm 0.02$ $0.96\pm 0.02$ $1.00\pm 0.00$ $0.33\pm 0.02$
Template $0.60\pm 0.14$ $0.17\pm 0.12$ $0.87\pm 0.10$ $0.32\pm 0.02$
NeMo + [Uncaptioned image] $150$ Verbatim $0.92\pm 0.02$ $0.96\pm 0.02$ $1.00\pm 0.00$ $0.32\pm 0.02$
Template $0.64\pm 0.15$ $0.16\pm 0.12$ $0.93\pm 0.06$ $0.32\pm 0.02$
Wanda + [Uncaptioned image] $1$ Verbatim $0.07\pm 0.02$ $0.07\pm 0.02$ $0.38\pm 0.10$ $0.21\pm 0.02$
Template $0.08\pm 0.03$ $0.08\pm 0.02$ $0.45\pm 0.12$ $0.22\pm 0.03$
Wanda + [Uncaptioned image] $10$ Verbatim $0.28\pm 0.10$ $0.28\pm 0.13$ $0.42\pm 0.07$ $0.29\pm 0.03$
Template $0.10\pm 0.04$ $0.21\pm 0.10$ $0.37\pm 0.09$ $0.28\pm 0.03$
Wanda + [Uncaptioned image] $25$ Verbatim $0.68\pm 0.10$ $0.70\pm 0.10$ $0.84\pm 0.13$ $0.31\pm 0.02$
Template $0.10\pm 0.05$ $0.68\pm 0.06$ $0.83\pm 0.08$ $0.31\pm 0.02$
Wanda + [Uncaptioned image] $50$ Verbatim $0.80\pm 0.06$ $0.84\pm 0.06$ $0.97\pm 0.02$ $0.32\pm 0.02$
Template $0.09\pm 0.05$ $0.78\pm 0.05$ $0.93\pm 0.04$ $0.31\pm 0.02$
Wanda + [Uncaptioned image] $100$ Verbatim $0.85\pm 0.05$ $0.87\pm 0.05$ $0.99\pm 0.01$ $0.32\pm 0.02$
Template $0.09\pm 0.06$ $0.82\pm 0.05$ $0.98\pm 0.01$ $0.32\pm 0.02$
Wanda + [Uncaptioned image] $150$ Verbatim $0.86\pm 0.04$ $0.87\pm 0.04$ $0.99\pm 0.00$ $0.32\pm 0.01$
Template $0.08\pm 0.07$ $0.82\pm 0.04$ $0.99\pm 0.01$ $0.32\pm 0.02$
E.4 Additional Hyperparameter Evaluation for Wanda
@@
Table 4: As shown, applying Wanda across all prompts is less effective at mitigating memorization compared to applying it individually per prompt. However, as discussed in Sec. E.7, applying Wanda per prompt and aggregating the found neurons over all 500 prompts comes at the high cost of reduced overall performance because of so many weights being pruned. In the setting with $10$ prompts, we randomly sample $10$ prompts across $5$ different seeds and report the average results. This setup proves less effective at mitigating memorization than using the full set of $500$ prompts.
Setting Memorization Type $\boldsymbol{\downarrow}\bm{\text{SSCD}_{\text{Orig}}}$ $\boldsymbol{\downarrow}\bm{\text{SSCD}_{\text{Gen}}}$ $\boldsymbol{\downarrow}\bm{D_{\text{SSCD}}}$ $\boldsymbol{\uparrow}\bm{A_{\text{CLIP}}}$
Wanda [3] per Prompt Verbatim $0.11\pm 0.03$ $0.12\pm 0.03$ $0.27\pm 0.06$ $0.32\pm 0.02$
Template $0.14\pm 0.04$ $0.13\pm 0.04$ $0.35\pm 0.10$ $0.32\pm 0.03$
Wanda [3] 10 Prompts Verbatim $0.22\pm 0.10$ $0.24\pm 0.11$ $0.41\pm 0.09$ $0.34\pm 0.02$
Template $0.19\pm 0.06$ $0.24\pm 0.13$ $0.42\pm 0.10$ $0.34\pm 0.03$
Wanda [3] all Prompts Verbatim $0.20\pm 0.08$ $0.21\pm 0.09$ $0.37\pm 0.07$ $0.34\pm 0.02$
Template $0.17\pm 0.05$ $0.18\pm 0.08$ $0.38\pm 0.09$ $0.34\pm 0.03$
Wanda per Prompt + [Uncaptioned image] Verbatim $0.69\pm 0.07$ $0.76\pm 0.06$ $0.91\pm 0.05$ $0.32\pm 0.02$
Template $0.10\pm 0.06$ $0.69\pm 0.06$ $0.86\pm 0.08$ $0.32\pm 0.02$
Wanda 10 Prompts + [Uncaptioned image] Verbatim $0.75\pm 0.06$ $0.81\pm 0.05$ $0.97\pm 0.02$ $0.32\pm 0.01$
Template $0.10\pm 0.06$ $0.72\pm 0.05$ $0.88\pm 0.06$ $0.32\pm 0.02$
Wanda all Prompts + [Uncaptioned image] Verbatim $0.80\pm 0.06$ $0.84\pm 0.06$ $0.97\pm 0.02$ $0.32\pm 0.02$
Template $0.09\pm 0.05$ $0.78\pm 0.05$ $0.93\pm 0.04$ $0.31\pm 0.02$
@@
Table 5: Even when applying Wanda [3] with a higher number of time steps it is still possible to break it using Dori.
Setting Number of Timesteps Memorization Type $\boldsymbol{\downarrow}\bm{\text{SSCD}_{\text{Orig}}}$ $\boldsymbol{\downarrow}\bm{\text{SSCD}_{\text{Gen}}}$ $\boldsymbol{\downarrow}\bm{D_{\text{SSCD}}}$ $\boldsymbol{\uparrow}\bm{A_{\text{CLIP}}}$
Wanda 1 Verbatim $0.22\pm 0.08$ $0.24\pm 0.09$ $0.38\pm 0.08$ $0.34\pm 0.02$
Template $0.18\pm 0.05$ $0.20\pm 0.08$ $0.39\pm 0.09$ $0.33\pm 0.02$
10 Verbatim $0.20\pm 0.08$ $0.21\pm 0.09$ $0.37\pm 0.07$ $0.34\pm 0.02$
Template $0.17\pm 0.05$ $0.18\pm 0.08$ $0.38\pm 0.09$ $0.34\pm 0.03$
20 Verbatim $0.20\pm 0.08$ $0.21\pm 0.07$ $0.39\pm 0.08$ $0.34\pm 0.03$
Template $0.17\pm 0.04$ $0.17\pm 0.06$ $0.39\pm 0.09$ $0.33\pm 0.03$
30 Verbatim $0.20\pm 0.08$ $0.20\pm 0.07$ $0.37\pm 0.07$ $0.34\pm 0.02$
Template $0.18\pm 0.05$ $0.17\pm 0.06$ $0.39\pm 0.10$ $0.33\pm 0.03$
40 Verbatim $0.20\pm 0.08$ $0.21\pm 0.07$ $0.38\pm 0.07$ $0.34\pm 0.02$
Template $0.18\pm 0.05$ $0.17\pm 0.07$ $0.39\pm 0.10$ $0.33\pm 0.03$
50 Verbatim $0.20\pm 0.07$ $0.20\pm 0.07$ $0.38\pm 0.07$ $0.34\pm 0.02$
Template $0.17\pm 0.04$ $0.17\pm 0.06$ $0.41\pm 0.11$ $0.33\pm 0.03$
Wanda + [Uncaptioned image] 1 Verbatim $0.77\pm 0.05$ $0.82\pm 0.05$ $0.97\pm 0.01$ $0.32\pm 0.01$
Template $0.10\pm 0.06$ $0.73\pm 0.05$ $0.88\pm 0.06$ $0.32\pm 0.02$
10 Verbatim $0.76\pm 0.05$ $0.82\pm 0.05$ $0.96\pm 0.02$ $0.32\pm 0.01$
Template $0.10\pm 0.06$ $0.72\pm 0.05$ $0.88\pm 0.06$ $0.32\pm 0.02$
20 Verbatim $0.76\pm 0.05$ $0.82\pm 0.05$ $0.96\pm 0.02$ $0.32\pm 0.01$
Template $0.11\pm 0.06$ $0.73\pm 0.05$ $0.87\pm 0.07$ $0.32\pm 0.02$
30 Verbatim $0.74\pm 0.05$ $0.80\pm 0.05$ $0.96\pm 0.02$ $0.32\pm 0.02$
Template $0.10\pm 0.05$ $0.73\pm 0.05$ $0.88\pm 0.07$ $0.32\pm 0.02$
40 Verbatim $0.74\pm 0.05$ $0.81\pm 0.05$ $0.96\pm 0.02$ $0.32\pm 0.01$
Template $0.10\pm 0.05$ $0.73\pm 0.05$ $0.88\pm 0.07$ $0.32\pm 0.02$
50 Verbatim $0.73\pm 0.06$ $0.79\pm 0.05$ $0.96\pm 0.02$ $0.32\pm 0.02$
Template $0.11\pm 0.05$ $0.72\pm 0.05$ $0.88\pm 0.06$ $0.32\pm 0.02$
E.5 Increasing Sparsity for Wanda
@@
Table 6: Applying Wanda [3] with higher sparsity does not change the fact that the method seems to only conceal memorization instead of completely removing it from the model. Increasing the sparsity also comes at the cost of reduced image quality, as the FID and the KID values suggest.
Setting Sparsity Memorization Type $\boldsymbol{\downarrow}\bm{\text{SSCD}_{\text{Orig}}}$ $\boldsymbol{\downarrow}\bm{\text{SSCD}_{\text{Gen}}}$ $\boldsymbol{\downarrow}\bm{D_{\text{SSCD}}}$ $\boldsymbol{\uparrow}\bm{A_{\text{CLIP}}}$ FID KID
Wanda 1% Verbatim $0.20\pm 0.08$ $0.21\pm 0.09$ $0.37\pm 0.07$ $0.34\pm 0.02$ $16.86$ $0.0067$
Template $0.17\pm 0.05$ $0.18\pm 0.08$ $0.38\pm 0.09$ $0.34\pm 0.03$ $17.51$ $0.0068$
2% Verbatim $0.19\pm 0.07$ $0.20\pm 0.07$ $0.36\pm 0.07$ $0.33\pm 0.02$ $18.17$ $0.0066$
Template $0.17\pm 0.05$ $0.16\pm 0.07$ $0.38\pm 0.08$ $0.33\pm 0.03$ $19.55$ $0.0073$
3% Verbatim $0.17\pm 0.07$ $0.17\pm 0.06$ $0.34\pm 0.06$ $0.33\pm 0.02$ $20.37$ $0.0075$
Template $0.16\pm 0.05$ $0.15\pm 0.06$ $0.38\pm 0.09$ $0.32\pm 0.03$ $22.40$ $0.0086$
4% Verbatim $0.17\pm 0.06$ $0.17\pm 0.05$ $0.34\pm 0.06$ $0.33\pm 0.02$ $23.07$ $0.0088$
Template $0.14\pm 0.05$ $0.14\pm 0.06$ $0.37\pm 0.09$ $0.32\pm 0.03$ $24.69$ $0.0097$
5% Verbatim $0.15\pm 0.05$ $0.16\pm 0.05$ $0.32\pm 0.05$ $0.32\pm 0.02$ $25.53$ $0.0102$
Template $0.13\pm 0.05$ $0.14\pm 0.05$ $0.39\pm 0.10$ $0.32\pm 0.03$ $26.61$ $0.0106$
10% Verbatim $0.12\pm 0.03$ $0.13\pm 0.04$ $0.33\pm 0.06$ $0.31\pm 0.02$ $37.34$ $0.0168$
Template $0.11\pm 0.04$ $0.13\pm 0.04$ $0.39\pm 0.10$ $0.30\pm 0.03$ $36.69$ $0.0166$
Wanda + [Uncaptioned image] 1% Verbatim $0.76\pm 0.06$ $0.82\pm 0.05$ $0.96\pm 0.02$ $0.32\pm 0.01$ $16.86$ $0.0067$
Template $0.10\pm 0.06$ $0.73\pm 0.05$ $0.88\pm 0.06$ $0.32\pm 0.02$ $17.51$ $0.0068$
2% Verbatim $0.71\pm 0.07$ $0.76\pm 0.06$ $0.90\pm 0.05$ $0.32\pm 0.02$ $18.17$ $0.0066$
Template $0.10\pm 0.05$ $0.67\pm 0.06$ $0.83\pm 0.08$ $0.32\pm 0.02$ $19.55$ $0.0073$
3% Verbatim $0.65\pm 0.08$ $0.73\pm 0.06$ $0.87\pm 0.06$ $0.31\pm 0.02$ $20.37$ $0.0075$
Template $0.10\pm 0.06$ $0.61\pm 0.08$ $0.76\pm 0.09$ $0.31\pm 0.02$ $22.40$ $0.0086$
4% Verbatim $0.62\pm 0.08$ $0.66\pm 0.08$ $0.81\pm 0.08$ $0.31\pm 0.02$ $23.07$ $0.0088$
Template $0.10\pm 0.05$ $0.58\pm 0.08$ $0.71\pm 0.10$ $0.31\pm 0.02$ $24.69$ $0.0097$
5% Verbatim $0.56\pm 0.10$ $0.62\pm 0.09$ $0.77\pm 0.10$ $0.31\pm 0.02$ $25.53$ $0.0102$
Template $0.10\pm 0.05$ $0.52\pm 0.10$ $0.66\pm 0.10$ $0.31\pm 0.02$ $26.61$ $0.0106$
10% Verbatim $0.40\pm 0.13$ $0.45\pm 0.14$ $0.67\pm 0.10$ $0.30\pm 0.02$ $37.34$ $0.0168$
Template $0.10\pm 0.04$ $0.35\pm 0.11$ $0.54\pm 0.07$ $0.30\pm 0.02$ $36.69$ $0.0166$
E.6 Iterative Application of NeMo

### v2

@@ H.3 Starting Embedding Optimization from Random Embeddings
[…]
NeMo + [Uncaptioned image] 11 Verbatim $07\pm 0.02 0.0$$06\pm 0.02$ $0.31\pm 0.06$ $21\pm 0.02 0.0$00.00
Template $0.10\pm 0.03$ $0.10\pm 0.03$ $0.26\pm 0.05$ $28\pm 0.02 0.0$00.00
NeMo + [Uncaptioned image] $10$ Verbatim $0.81\pm 0.06$ $0.89\pm 0.05$ $0.99\pm 0.01$ $0.32\pm 0.01$ $0.79$
Template $0.42\pm 0.13$ $0.21\pm 0.15$ $0.72\pm 0.14$ $0.32\pm 0.02$ $0.14$
NeMo + [Uncaptioned image] $25$ Verbatim $0.88\pm 0.04$ $0.95\pm 0.03$ $1.00\pm 0.00$ $0.32\pm 0.01$ $0.94$
Template $0.50\pm 0.13$ $0.20\pm 0.15$ $0.75\pm 0.11$ $0.32\pm 0.02$ $0.13$
NeMo + [Uncaptioned image] $50$ Verbatim $0.91\pm 0.03$ $0.97\pm 0.02$ $1.00\pm 0.00$ $0.33\pm 0.02$ $0.99$
Template $0.55\pm 0.12$ $0.18\pm 0.12$ $0.79\pm 0.10$ $0.32\pm 0.02$ $0.12$
NeMo + [Uncaptioned image] $100$ Verbatim $0.93\pm 0.02$ $0.96\pm 0.02$ $1.00\pm 0.00$ $0.33\pm 0.02$ $1.00$
Template $0.60\pm 0.14$ $0.17\pm 0.12$ $0.87\pm 0.10$ $0.32\pm 0.02$ $0.10$
NeMo + [Uncaptioned image] $150$ Verbatim $0.92\pm 0.02$ $0.96\pm 0.02$ $1.00\pm 0.00$ $0.32\pm 0.02$ $1.00$
Template $0.64\pm 0.15$ $0.16\pm 0.12$ $0.93\pm 0.06$ $32\pm 0.02 0.0$70.07
Wanda + [Uncaptioned image] 11 Verbatim $07\pm 0.02 0.0$$07\pm 0.02$ $0.38\pm 0.10$ $21\pm 0.02 0.0$00.00
Template $07\pm 0.02 0.0$$07\pm 0.02$ $0.43\pm 0.12$ $22\pm 0.02 0.0$00.00
Wanda + [Uncaptioned image] $10$ Verbatim $0.28\pm 0.10$ $0.28\pm 0.13$ $0.42\pm 0.07$ $29\pm 0.03 0.0$20.02
Template $0.27\pm 0.09$ $0.12\pm 0.05$ $0.47\pm 0.16$ $29\pm 0.02 0.0$00.00
Wanda + [Uncaptioned image] $25$ Verbatim $0.68\pm 0.10$ $0.70\pm 0.10$ $0.84\pm 0.13$ $0.31\pm 0.02$ $0.37$
Template $0.46\pm 0.11$ $0.15\pm 0.08$ $0.68\pm 0.09$ $0.31\pm 0.02$ $0.10$
Wanda + [Uncaptioned image] $50$ Verbatim $0.80\pm 0.06$ $0.84\pm 0.06$ $0.97\pm 0.02$ $0.32\pm 0.02$ $0.78$
Template $0.54\pm 0.12$ $0.14\pm 0.08$ $0.76\pm 0.10$ $0.31\pm 0.02$ $0.26$
Wanda + [Uncaptioned image] $100$ Verbatim $0.85\pm 0.05$ $0.87\pm 0.05$ $0.99\pm 0.01$ $0.32\pm 0.02$ $0.89$
Template $0.60\pm 0.15$ $0.15\pm 0.10$ $0.84\pm 0.10$ $0.32\pm 0.02$ $0.37$
Wanda + [Uncaptioned image] $150$ Verbatim $0.86\pm 0.04$ $0.87\pm 0.04$ $0.99\pm 0.00$ $0.32\pm 0.01$ $0.92$
Template $0.64\pm 0.16$ $0.16\pm 0.10$ $0.90\pm 0.08$ $0.31\pm 0.03$ $0.44$
H.4 Additional Hyperparameter Evaluation for Wanda
@@
Table 6: As shown, applying Wanda across all prompts is less effective at mitigating memorization compared to applying it individually per prompt. However, as discussed in Section˜H.7, applying Wanda per prompt and aggregating the found neurons over all 500 prompts comes at the high cost of reduced overall performance because of so many weights being pruned. In the setting with $10$ prompts, we randomly sample $10$ prompts across $5$ different seeds and report the average results. This setup proves less effective at mitigating memorization than using the full set of $500$ prompts.
[…]
Wanda (Chavhan et al., 2024) per Prompt Verbatim $0.11\pm 0.03$ $0.12\pm 0.03$ $0.27\pm 0.06$ $32\pm 0.02 0.0$00.00
Template $0.14\pm 0.04$ $0.13\pm 0.04$ $0.35\pm 0.10$ $32\pm 0.03 0.0$00.00
Wanda (Chavhan et al., 2024) 10 Prompts Verbatim $0.22\pm 0.10$ $0.24\pm 0.11$ $0.41\pm 0.09$ $34\pm 0.02 0.0$10.01
Template $0.19\pm 0.06$ $0.24\pm 0.13$ $0.42\pm 0.10$ $34\pm 0.03 0.0$10.01
Wanda (Chavhan et al., 2024) all Prompts Verbatim $0.20\pm 0.08$ $0.21\pm 0.09$ $0.37\pm 0.07$ $34\pm 0.02 0.0$00.00
Template $0.17\pm 0.05$ $0.18\pm 0.08$ $0.38\pm 0.09$ $34\pm 0.03 0.0$00.00
Wanda per Prompt + [Uncaptioned image] Verbatim $0.69\pm 0.07$ $0.76\pm 0.06$ $0.91\pm 0.05$ $0.32\pm 0.02$ $0.46$
Template $0.52\pm 0.10$ $0.17\pm 0.10$ $0.75\pm 0.07$ $0.32\pm 0.02$ $0.15$
Wanda 10 Prompts + [Uncaptioned image] Verbatim $0.75\pm 0.06$ $0.81\pm 0.05$ $0.97\pm 0.02$ $0.32\pm 0.01$ $0.71$
Template $0.50\pm 0.11$ $0.16\pm 0.09$ $0.74\pm 0.09$ $0.32\pm 0.02$ $0.15$
Wanda all Prompts + [Uncaptioned image] Verbatim $0.76\pm 0.05$ $0.82\pm 0.05$ $0.96\pm 0.02$ $0.32\pm 0.01$ $0.73$
Template $0.51\pm 0.11$ $0.16\pm 0.09$ $0.75\pm 0.08$ $0.32\pm 0.02$ $0.16$
@@
Table 7: Even when applying Wanda Chavhan et al. (2024) with a higher number of time steps it is still possible to break it using Dori.
[…]
Wanda 1 Verbatim $0.22\pm 0.08$ $0.24\pm 0.09$ $0.38\pm 0.08$ $34\pm 0.02 0.0$10.01
Template $0.18\pm 0.05$ $0.20\pm 0.08$ $9\pm 0.09 0.33±0$$3\pm 0.02$ $0.00$
10 Verbatim $0.20\pm 0.08$ $0.21\pm 0.09$ $0.37\pm 0.07$ $34\pm 0.02 0.0$00.00
Template $0.17\pm 0.05$ $0.18\pm 0.08$ $0.38\pm 0.09$ $34\pm 0.03 0.0$00.00
20 Verbatim $0.20\pm 0.08$ $0.21\pm 0.07$ $0.39\pm 0.08$ $34\pm 0.03 0.0$00.00
Template $0.17\pm 0.04$ $0.17\pm 0.06$ $9\pm 0.09 0.33±0$$3\pm 0.03$ $0.00$
30 Verbatim $0.20\pm 0.08$ $0.20\pm 0.07$ $0.37\pm 0.07$ $34\pm 0.02 0.0$00.00
Template $0.18\pm 0.05$ $0.17\pm 0.06$ $9\pm 0.10 0.33±0$$3\pm 0.03$ $0.00$
40 Verbatim $0.20\pm 0.08$ $0.21\pm 0.07$ $0.38\pm 0.07$ $34\pm 0.02 0.0$00.00
Template $0.18\pm 0.05$ $0.17\pm 0.07$ $9\pm 0.10 0.33±0$$3\pm 0.03$ $0.00$
50 Verbatim $0.20\pm 0.07$ $0.20\pm 0.07$ $0.38\pm 0.07$ $34\pm 0.02 0.0$00.00
Template $0.17\pm 0.04$ $0.17\pm 0.06$ $0.41\pm 0.11$ $33\pm 0.03 0.0$00.00
Wanda + [Uncaptioned image] 1 Verbatim $0.77\pm 0.05$ $0.82\pm 0.05$ $0.97\pm 0.01$ $0.32\pm 0.01$ $0.79$
Template $0.51\pm 0.11$ $0.16\pm 0.09$ $0.75\pm 0.09$ $0.32\pm 0.02$ $0.18$
10 Verbatim $0.76\pm 0.05$ $0.82\pm 0.05$ $0.96\pm 0.02$ $0.32\pm 0.01$ $0.73$
Template $0.52\pm 0.11$ $0.16\pm 0.09$ $0.75\pm 0.09$ $0.32\pm 0.02$ $0.15$
20 Verbatim $0.76\pm 0.05$ $0.82\pm 0.05$ $0.96\pm 0.02$ $0.32\pm 0.01$ $0.72$
Template $0.52\pm 0.11$ $0.16\pm 0.09$ $0.75\pm 0.09$ $0.32\pm 0.02$ $0.17$
30 Verbatim $0.74\pm 0.05$ $0.80\pm 0.05$ $0.96\pm 0.02$ $0.32\pm 0.02$ $0.68$
Template $0.51\pm 0.11$ $0.16\pm 0.09$ $0.75\pm 0.09$ $0.32\pm 0.02$ $0.14$
40 Verbatim $0.74\pm 0.05$ $0.81\pm 0.05$ $0.96\pm 0.02$ $0.32\pm 0.01$ $0.69$
Template $0.50\pm 0.12$ $0.16\pm 0.09$ $0.74\pm 0.09$ $0.32\pm 0.02$ $0.18$
50 Verbatim $0.73\pm 0.06$ $0.79\pm 0.05$ $0.96\pm 0.02$ $0.32\pm 0.02$ $0.59$
Template $0.49\pm 0.12$ $0.16\pm 0.09$ $0.74\pm 0.09$ $0.32\pm 0.02$ $0.14$
H.5 Increasing Sparsity for Wanda
@@
Table 8: Applying Wanda Chavhan et al. (2024) with higher sparsity does not change the fact that the method seems to only conceal memorization instead of completely removing it from the model. Increasing the sparsity also comes at the cost of reduced image quality, as the FID and the KID values suggest.
[…]
Wanda 1% Verbatim $0.20\pm 0.08$ $0.21\pm 0.09$ $0.37\pm 0.07$ $34\pm 0.02 0.0$00.00 $16.86$ $0.0067$
Template $0.17\pm 0.05$ $0.18\pm 0.08$ $0.38\pm 0.09$ $34\pm 0.03 0.0$00.00 $17.51$ $0.0068$
2% Verbatim $0.19\pm 0.07$ $0.20\pm 0.07$ $6\pm 0.07 0.33±0$$3\pm 0.02$ $0.00$ $18.17$ $0.0066$
Template $0.17\pm 0.05$ $0.16\pm 0.07$ $8\pm 0.08 0.33±0$$3\pm 0.03$ $0.00$ $19.55$ $0.0073$
3% Verbatim $0.17\pm 0.07$ $0.17\pm 0.06$ $4\pm 0.06 0.33±0$$3\pm 0.02$ $0.00$ $20.37$ $0.0075$
Template $0.16\pm 0.05$ $0.15\pm 0.06$ $0.38\pm 0.09$ $32\pm 0.03 0.0$00.00 $22.40$ $0.0086$
4% Verbatim $0.17\pm 0.06$ $0.17\pm 0.05$ $4\pm 0.06 0.33±0$$3\pm 0.02$ $0.00$ $23.07$ $0.0088$
Template $0.14\pm 0.05$ $0.14\pm 0.06$ $0.37\pm 0.09$ $32\pm 0.03 0.0$00.00 $24.69$ $0.0097$
5% Verbatim $0.15\pm 0.05$ $0.16\pm 0.05$ $0.32\pm 0.05$ $32\pm 0.02 0.0$00.00 $25.53$ $0.0102$
Template $0.13\pm 0.05$ $0.14\pm 0.05$ $0.39\pm 0.10$ $32\pm 0.03 0.0$00.00 $26.61$ $0.0106$
10% Verbatim $0.12\pm 0.03$ $0.13\pm 0.04$ $0.33\pm 0.06$ $31\pm 0.02 0.0$00.00 $37.34$ $0.0168$
Template $0.11\pm 0.04$ $0.13\pm 0.04$ $0.39\pm 0.10$ $30\pm 0.03 0.0$00.00 $36.69$ $0.0166$
Wanda + [Uncaptioned image] 1% Verbatim $0.76\pm 0.06$ $0.82\pm 0.05$ $0.96\pm 0.02$ $0.32\pm 0.01$ $0.73$ $16.86$ $0.0067$
Template $0.51\pm 0.12$ $0.16\pm 0.09$ $0.75\pm 0.09$ $0.32\pm 0.02$ $0.16$ $17.51$ $0.0068$
2% Verbatim $0.71\pm 0.07$ $0.76\pm 0.06$ $0.90\pm 0.05$ $0.32\pm 0.02$ $0.54$ $18.17$ $0.0066$
Template $0.45\pm 0.11$ $0.15\pm 0.08$ $0.71\pm 0.08$ $31\pm 0.03 0.0$90.09 $19.55$ $0.0073$
3% Verbatim $0.65\pm 0.08$ $0.73\pm 0.06$ $0.87\pm 0.06$ $0.31\pm 0.02$ $0.32$ $20.37$ $0.0075$
Template $0.39\pm 0.10$ $0.13\pm 0.07$ $0.70\pm 0.08$ $31\pm 0.03 0.0$40.04 $22.40$ $0.0086$
4% Verbatim $2\pm 0.08 0.66±0$$6\pm 0.08$ $0.81\pm 0.08$ $0.31\pm 0.02$ $0.17$ $23.07$ $0.0088$
Template $0.35\pm 0.12$ $0.14\pm 0.07$ $0.68\pm 0.08$ $31\pm 0.03 0.0$30.03 $24.69$ $0.0097$
5% Verbatim $0.56\pm 0.10$ $0.62\pm 0.09$ $0.77\pm 0.10$ $31\pm 0.02 0.0$80.08 $25.53$ $0.0102$
Template $0.29\pm 0.10$ $0.13\pm 0.07$ $0.68\pm 0.09$ $30\pm 0.04 0.0$20.02 $26.61$ $0.0106$
10% Verbatim $0.40\pm 0.13$ $0.45\pm 0.14$ $0.67\pm 0.10$ $30\pm 0.02 0.0$10.01 $37.34$ $0.0168$
Template $0.22\pm 0.08$ $0.12\pm 0.06$ $0.63\pm 0.10$ $28\pm 0.04 0.0$10.01 $36.69$ $0.0166$
H.6 Iterative Application of NeMo